# Towards Automated Identification of Violation Symptoms of Architecture Erosion

### v4

@@ 5.1.1. RQ1: Identifying Violation Symptoms
Our trained ML/DL classifiers can identify violation symptoms in code review comments. The results of RQ1.1 indicate that certain factors such as the type of classifier, the pre-trained word embeddings, and hyperparameters influence the performance of classifiers. Specifically, we found that the SVM classifier with word2vec outperforms other ML and DL-based classifiers. We also found that DL classifiers do not always have better performance than ML classifiers; we note that training DL classifiers takes much longer time than training ML classifiers, as DL classifiers have much more complex network architecture and more parameters. In recent years, some studies that leveraged DL-based techniques for software engineering tasks have achieved better performance that all other state-of-the-art techniques (e.g., ML-based techniques) (Yang2022pmse, ). However, the achieved results imply that ML classifiers might be enough for certain text classification tasks with small-size datasets, especially as the trained DL classifiers do not show performance improvement in such a scenario. This finding is also observed in previous studies (Liu2018nmt, ; Fu2017eoh, ): for certain software engineering tasks, simple ML-based approaches are capable of achieving equal or even better performance than DL-based approaches within less time. Thus, we advise researchers to not blindly apply DL-based techniques and consider simple approaches first.
Finding 1: ML-based approaches have sufficient performance for many software engineering tasks, such as identification of architecture violations in code reviews.
@@
Besides, the results of RQ1.3 show that 200-dimensional pre-trained word embedding models outperform 100 and 300-dimensional models. Our results are slightly different from previous studies (e.g., (Ren2019nndSATD, ; Li2022SATD, )) where the 300-dimensional word embedding model has better performance. One possible reason is that the previous studies only measured the average results of classifiers with different dimensional word embedding models, and the average results might be influenced by the individual results. Our results present and compare the classification performance of each ML/DL classifier, which could be more accurate.
@@ 5.1.2. RQ2: Improving Performance with Voting
Evaluating the performance of an ensemble classifier offers insight into the performance improvement compared with individual classifiers. For voting-based approaches, an accurate output is attained only when a majority of the constituent models generate the correct output (hansen1990nne, ). Our results provide further empirical evidence for the efficacy of the majority voting strategy regarding improving the performance of text classification. This technique effectively combines the predictions of multiple classifiers to arrive at a final decision, resulting in improved overall classification performance. Nevertheless, we also observed that such performance improvement is not obvious in TextCNN in our case. One possible reason is that we only get an ensemble DL classifier based on three individual TextCNN models, and that causes the voting results to be particularly susceptible to the influence of single results. We argue that the voting strategy remains valuable to minimize potential prediction errors if there are more outputs for voting.
Finding 3: Compared with individual classifiers for identifying architecture violations from code review comments, ensemble learning is an effective way for improving classification performance. In most cases, ensemble classifiers that utilize the voting strategy can enhance all classification performance metrics.
@@ 5.1.3. RQ3: Validation in Practice
The existence of architectural violations can be challenging, as they can impede practitioners’ understanding of architecture during the development process. This particularly influences novice developers who lack practical familiarity with architecture. The results of RQ3 indicate a predominantly positive feedback, confirming the usefulness of our trained ML/DL classifiers in potentially helping developers during maintenance and evolution. Through our analysis of the received responses to the statements (e.g., Statements 3 and 5), we found that developers expressed a significant interest in utilizing automated checking approaches, tools, or plug-ins to identify violation symptoms during code review, recognizing their potential to enhance the productivity of dealing with architectural violations through code review. The survey and interview results indicate that developers hold positive attitudes towards such kind of tools. Moreover, such automated techniques regarding identifying potential architecture violations could be supplementary means to minimize architecture violations and improve software quality during code review. For example, such techniques could provide code reviewers with warnings that architecture violations need to be refactored, based on the review discussions. Furthermore, they could help code reviewers prioritize such refactorings based on various criteria (e.g., the importance of involved components, the number of violation symptoms in those components, the seniority of code reviewers). Such a prioritization could help to fix violation symptoms that might otherwise have been ignored.
@@
Moreover, according to the feedback from respondents, certain practitioners might hold a conservative attitude toward ML and DL-based classifiers for identifying violation symptoms, which could be observed from the “Neutral” results in Figure 4, especially for Statement 2. One possible reason could be that such identification techniques merely provide warnings about potential issues, rather than offering refactoring suggestions to effectively address those violation symptoms. Another possible reason is that they might be skeptical about the performance of the classifiers (false positives or negatives), casting doubt about the reliability of identified architectural violations. Nevertheless, no classifier can ever achieve 100 % accuracy in classification tasks, and there is always a trade-off between precision and recall. This disclaimer should be made clear when showing classification tasks to practitioners. Another potential reason could be that certain identified review comments may be part of the architectural design discussions during the development process, rather than the final architectural decisions. Such concerns are beyond the target of our research, as the trained classifiers can only provide judgments based on the collected violations, instead of exploring the evolution of architecture violations.
Finding 4: Despite a few concerns about the accuracy of the identified architecture violations, the majority of positive feedback from participants indicates that practitioners are receptive to the use of automated techniques for identifying architecture violations during code review.
@@ 5.2. Implications
Choose appropriate classifiers. An ensemble classifier is a collection of classifiers that work together to classify instances by combining their individual outputs through voting. Researchers can consider exploring different selections of classifiers to further enhance the performance of ensemble classifiers. For example, to some extent, assigning varying weights to individual classifiers might partially mitigate the impact of biased classifiers on voting results.
Suggestion 2: It is recommended for researchers to explore various combinations of classifiers and assign different weights to individual classifiers to further enhance the classification performance of ensemble classifiers for identifying architecture violations.
Take time investment into consideration. Even though our study did not specifically investigate the time efficiency of classifiers’ training, we observed that the TextCNN classifiers exhibited a long training time due to a large number of parameters. In comparison to traditional ML classifiers, DL classifiers require a longer training time, yet the results did not demonstrate a significant performance improvement. Therefore, when considering the practical implementation of classifiers, it is important to take into account the cost-benefit ratio of “Performance/Time”. In situations where there is a limited dataset, traditional machine learning techniques may suffice in achieving the desired goals. Further research may explore additional methods for balancing the trade-off between performance and training time for classifiers.
Suggestion 3: For certain software engineering tasks, constructing simple classifiers (e.g., individual or ensemble classifiers) can yield a superior return on investment compared to training complex DL-based classifiers, particularly when time investment needs to be taken into account.
Optimize code review practice. In the future, researchers and practitioners can explore the development of plug-ins for code review platforms (e.g., Gerrit) that provide warnings of violations to developers or even suggestions to fix those violations. Such plug-ins can be developed utilizing historical code review data or harnessing the existing large language models through fine-tuning. These plug-ins have the potential to assist developers and maintainers in conducting comprehensive examinations of preexisting architectural violations, and may even serve as reminders of their presence. Additionally, implementing the trained classifiers (e.g., as an extension) into the existing tools (or toolsets) is also worthy of consideration in practice. For example, code review platforms can integrate such classifiers to empower the automated code review process. Such tools may provide more insights and help novices or maintainers troubleshoot architecturally-relevant violation issues.
Suggestion 4: Researchers and practitioners can consider developing plug-ins for code review platforms (e.g., Gerrit) to assist developers in identifying potential architecture violations, or integrating the trained classifiers into existing tools to optimize the automated code review process.
@@ 6. Threats to Validity
In this section, we discuss the potential threats to the validity related to our study and the adopted measures mitigating these threats by following the guidelines proposed by Wohlin et al. (Wohlin2012ESE, ).
@@ 6.1. Construct validity
Construct validity concerns the connection between operational measures and the studied subjects. The first potential threat to construct validity is the suitability of our selected ML/DL algorithms, feature extraction techniques, and metrics. As mentioned by Peter et al. (peters2017text, ), it is not possible to use an exhaustive approach to test all learning models. We selected five common ML algorithms, one DL algorithm, and three popular pre-trained word embedding models, as they are all widely used to conduct text classification. We admit that using different algorithms and pre-trained word embedding models may generate varying results, despite their popularity. We cannot completely mitigate this threat. In addition, the four metrics (recall, precision, F1-score, and accuracy) are widely used to evaluate the performance of various automatic software engineering techniques (see Section 7.3). Therefore, the metrics pose little threat to the construct validity.
Another potential threat comes from the convenience sampling method of the survey and interview, as individuals with negative views towards the research might be less inclined to participate in our survey and interview. This is a compromise inherent to the convenience sampling method, and it implies that this threat cannot be completely mitigated. Nevertheless, to improve the response rate and collect as many opinions as possible from practitioners, we sent two rounds of reminder emails to the involved practitioners. It is also possible that some survey respondents did not understand the statements well. To reduce this threat to construct validity, we conducted a pilot survey for refining the survey design. Besides, we provided an “I Don’t Know” option in our survey. Moreover, we also interviewed six participants to collect in-depth opinions besides the six defined statements in the survey. To mitigate this threat, we offered an optional open-ended question that allows respondents to freely share their thoughts with us.
@@ 6.2. External validity
External validity pertains to the generalizability of our experiment results and findings. We trained the ML and DL-based classifiers with our collected dataset regarding architecture violations. Selecting different datasets and experimental settings might influence the generalizability of (part of) our findings. We acknowledge that the chosen open-source projects are not representative of all projects from the OpenStack and Qt platforms, and our findings may not be generalized to diverse projects, particularly those from different domains such as closed-source projects or projects on other platforms like Apache. Furthermore, the participants of the survey and interviews are contributors to the selected projects, therefore, the findings of RQ3 might not be generalized to other projects. To support further generalizability, we have made our dataset (onlinepackage_TSE, ) available to enable other researchers and practitioners to replicate, reuse, and extend our study.
@@ 6.3. Reliability
Reliability reflects the replicability of a study regarding yielding the same or similar results when other researchers reproduce this study. One potential threat to reliability comes from the experimental settings (e.g., input data and hyperparameters for training classifiers) used in this work. We used the dataset from our previous study (Li2023wvs, ) that had publicly shared the labeled data and corresponding scripts for collecting data. Besides, we also shared the source code in our replication package (onlinepackage_TSE, ) to facilitate reproduction and validation. Another potential threat might be from the survey and interview design. To mitigate the threat, we provided the templates of our survey, semi-structured interviews, and customized emails in our replication package (onlinepackage_TSE, ) to encourage further research in this area.
@@ 7. Related Work
In this section, we briefly discuss the current state of research in three associated areas of our study. We first describe related work on code review comments in Section 7.1; we then introduce previous studies on architecture violations in Section 7.2; we finally present the benefits of utilizing ML and DL techniques to analyze artifacts in software development in Section 7.3.
@@ 7.1. Code Review Comments
While the aforementioned studies investigated various aspects of code review regarding development and maintenance (e.g., decisions, design degradation, and SATD), there are no studies that investigated the automated identification of violation symptoms in code review comments; we decided to fill this gap and explore the application of ML/DL techniques to specifically identify violation-related issues (e.g., violation symptoms) from code review comments.
@@ 7.2. Architecture Violations
Besides, Terra et al. (Terra2015rsr, ) reported their experience in fixing architectural violations. They proposed a recommendation system ArchFix that provides refactoring guidelines for developers and maintainers to repair architectural violations in the module architecture view of object-oriented systems. Their results show that their approach can provide correct recommendations for 75% and 79% of the detected architecture violations on two selected systems, respectively. Maffort et al. (Maffort2016mav, ) proposed an approach to check architecture conformance for detecting architecture violations based on defined heuristics. They claimed that their approach relies on the defined heuristic rules and can rapidly raise architecture violation warnings.
Different from the abovementioned studies focusing on detecting architecture violations in source code, our work explores the possibility of automated identification of the violation symptoms from code review comments.
7.3. Analyzing Software Repositories with Machine Learning and Deep Learning

### v5

@@ 5.1.1. RQ1: Identifying Violation Symptoms
Our trained ML/DL classifiers can identify violation symptoms in code review comments. The results of RQ1.1 indicate that certain factors such as the type of classifier, the pre-trained word embeddings, and hyperparameters influence the performance of classifiers. Specifically, we found that the SVM classifier with word2vec outperforms other ML/DL-based classifiers. We also found that DL classifiers do not always have better performance than ML classifiers, and the outcome may stem from factors such as model complexity, dataset size, and hyperparameter settings; we note that training DL classifiers takes much longer time than training ML classifiers, as DL classifiers have much more complex network architecture and more parameters. In recent years, some studies that leveraged DL-based techniques for SE tasks have achieved better performance that all other state-of-the-art techniques (e.g., ML-based techniques) (Yang2022pmse, ). However, the achieved results imply that ML classifiers might be enough for certain text classification tasks with small-size datasets, especially as the trained DL classifiers do not show performance improvement in such a scenario. This finding is also observed in previous studies (Liu2018nmt, ; Fu2017eoh, ): for certain SE tasks, simple ML-based approaches are capable of achieving equal or even better performance than DL-based approaches within less time. Thus, we advise researchers to not blindly apply DL-based techniques and consider simple approaches first.
Finding 1: Traditional ML-based approaches have sufficient performance for many SE tasks, such as identification of architecture violations in code reviews.
@@
Besides, the results of RQ1.3 show that 200-dimensional pre-trained word embedding models outperform 100 and 300-dimensional models. The possible reason is that 200-dimensional embeddings strike a balance between expressiveness and robustness for our dataset size: lower-dimensional (100d) embeddings under-represent nuanced lexical patterns, whereas higher-dimensional (300d) embeddings introduce excess parameters, increasing the risk of overfitting. Our results are slightly different from previous studies (e.g., (Ren2019nndSATD, ; Li2022SATD, )) where the 300-dimensional word embedding model has better performance. One possible reason is that the previous studies only measured the average results of classifiers with different dimensional word embedding models, and the average results might be influenced by the individual results. Our results present and compare the classification performance of each ML/DL classifier, which could be more accurate.
@@ 5.1.2. RQ2: Improving Performance with Voting
Evaluating the performance of an ensemble classifier offers insight into the performance gains compared with individual classifiers. For voting-based approaches, an accurate output is attained only when a majority of the constituent models generate the correct output (hansen1990nne, ). Our results provide further empirical evidence for the efficacy of the majority voting strategy on improving the text classification performance. This technique effectively combines the predictions of multiple classifiers to reach a final decision, leading to improved overall classification performance. Nevertheless, we observed that such performance improvement is not obvious in TextCNN in our case. One possible reason is that our DL classifiers (all TextCNN variants) and the modest dataset size prevented the majority‑vote ensemble from yielding significant gains. We argue that the voting strategy remains valuable to minimize potential prediction errors if there are more outputs for voting.
Finding 3: Compared with individual ML/DL-based classifiers for identifying architecture violations from code review comments, ensemble learning is an effective way for improving classification performance. In most cases, ensemble classifiers that utilize the voting strategy can enhance all classification performance metrics.
@@ 5.1.3. RQ3: Validation in Practice
The existence of architectural violations can be challenging, as they can impede practitioners’ understanding of architecture. This particularly influences novice developers who lack practical familiarity with architecture. The practitioners we surveyed and interviewed recognized the identified issues in code snippets, as they had previously noted them during code review. However, they may not have realized these issues were violation symptoms of architecture erosion. Our approach helps practitioners realize the existence of violation symptoms of architecture erosion. The results of RQ3 indicate a predominantly positive feedback, confirming the usefulness of our trained ML/DL classifiers in potentially helping developers during maintenance and evolution. Through analyzing the received responses to the statements (especially Statements 3, 4, and 5), we found that developers convey interest in utilizing automated approaches, tools, or plug-ins to identify violation symptoms during code review, recognizing their potential to enhance the productivity of dealing with architectural violations. Overall, the survey and interview results indicate that developers hold positive attitudes towards such kinds of tools. To some extent, their perception might be influenced by the use of balanced datasets in classifier evaluation. Future work could focus on deploying the classifiers in more real-world scenarios, where data distributions may differ and classifiers’ performance could vary, potentially improving or declining compared to our current results. Moreover, automated techniques regarding identifying potential architecture violations could be supplementary means to minimize architecture violations and improve code quality during code review. For example, such techniques could provide code reviewers with warnings that architecture violations need to be refactored, based on the review discussions. Furthermore, according to the participants’ feedback, automated techniques could help code reviewers prioritize such refactorings based on various criteria (e.g., the importance of involved components, the number of violation symptoms in those components, and the seniority of code reviewers). Such a prioritization could help to fix violation symptoms that might otherwise have been ignored.
@@
Moreover, according to the respondents’ feedback, certain practitioners might hold a conservative attitude toward ML/DL-based classifiers for identifying violation symptoms, which could be observed from the “Neutral” results in Figure 4, especially for Statement 2. One reason is that they might be skeptical about the performance of the classifiers (false positives or negatives), casting doubt about the reliability of identified violation symptoms. Nevertheless, no classifier can ever achieve 100% accuracy in classification tasks, and there is always a trade-off between precision and recall. This disclaimer should be made clear when showing classification tasks to practitioners. Another potential reason could be that certain identified review comments may be part of the architectural design discussions during the development process, rather than the final architectural decisions. This is particularly common in agile development, where architectures evolve over time and certain decisions are continuously discussed and refined. Such concerns are beyond the target of our research, as the trained classifiers can only provide judgments based on the collected violations, instead of exploring the evolution of architecture violations.
Finding 4: Despite a few concerns about the accuracy of the identified violation symptoms, the vast majority reacted favorably, indicating strong practitioner support for automated identification of violation symptoms during code review.
5.1.4. RQ4: Comparison: ML/DL-based vs. LLM-based Classifiers
The results of RQ4 clearly demonstrate that LLMs when configured as classifiers via prompt engineering, substantially outperform traditional ML/DL-based approaches on the violation symptom identification task. Across all metrics, the average performance of LLM-based classifiers exceeds that of both ML/DL-based classifiers. This suggests that LLMs’ deep contextual understanding and few-shot adaptability confer distinct advantages for text classification tasks.
Among the three LLM-based classifiers, the results show that different LLMs vary in performance on the automated identification of violation symptoms. The possible reason could be that different LLMs draw on distinct training corpora and pre-training goals, which in turn shape their strengths and weaknesses across SE tasks, such as program repair, code generation. Consequently, practitioners should align task requirements (e.g., code understanding, text classification) with each LLM’s underlying training focus to select suitable LLMs.
Furthermore, ensemble strategy either across models (see Table 10) or across shot settings for an individual model (see Table 10) yields modest but consistent improvements over average results of various shot settings from Table 8. However, ensemble classifiers generally do not outperform the best-performing individual LLMs in Precision or Recall, underscoring that an ensemble’s strength lies in balancing errors rather than maximizing peak performance capabilities. Notably, employing a voting strategy cannot significantly improve the performance of ensemble LLM-based classifiers. A plausible explanation could be that the individual LLM-based classifiers already achieve very good performance, leaving limited room for further enhancement through ensembling. In fact, the ensemble technique may even degrade performance, potentially due to the amplification of inherent issues such as LLM-induced hallucination (GPT4, ). This phenomenon closely resembles the Law of Diminishing Marginal Utility in economics, whereby additional inputs yield progressively smaller benefits beyond a certain threshold (jevons2013PE, ).
Finding 5: LLM-based classifiers typically outperform traditional ML/DL-based classifiers for identifying violation symptoms. However, ensembling multiple LLM-based classifiers does not yield further performance gains compared to individual LLM-based classifiers, potentially due to the compounding effect of hallucinations inherent in LLMs.
@@ 5.2. Implications
Choose appropriate classifiers. For traditional ML/DL-based classifiers, an ensemble classifier is a collection of classifiers that work together to classify instances by combining their individual outputs through voting. Researchers can consider exploring different selections of classifiers to further enhance the performance of ensemble classifiers. For example, to some extent, assigning varying weights to individual classifiers might partially mitigate the impact of biased classifiers on voting results. While ensemble classifiers based on LLMs require more time and resources, their performance often does not exceed that of certain top-performing individual LLM-based classifiers, particularly when configured with optimal few-shot settings.
Suggestion 2: When constructing traditional ML/DL-based classifiers, combining classifiers and assigning different weights to individual classifiers are recommended to further enhance the classification performance of ensemble classifiers for identifying architecture violations. In contrast, for LLM-based classifiers, choosing a suitable few-shot setting is recommended over ensemble techniques.
Optimize code review practice. In the future, researchers and practitioners can explore the development of plug-ins for code review platforms (e.g., Gerrit) that provide warnings of violations to developers or even suggestions to fix those violations. Such plug-ins can be developed utilizing historical code review data or harnessing the existing LLMs through fine-tuning. These plug-ins have the potential to assist developers and maintainers in conducting comprehensive examinations of preexisting architectural violations, and may even serve as reminders of their presence. Additionally, implementing the trained classifiers (e.g., as an extension) into the existing tools (or toolsets) is also worthy of consideration in practice. For example, code review platforms can integrate such classifiers to empower the automated code review process. Such tools may provide more insights and help novices or maintainers troubleshoot architecturally relevant violation issues.
Suggestion 3: Researchers and practitioners can consider integrating classifiers as plug-ins for code review platforms (e.g., Gerrit) to assist developers in identifying potential architecture violations, or integrating customized classifiers into existing tools to optimize the automated code review process.
@@ 6. Threats to Validity
In this section, we discuss the potential threats to the validity related to our study and the adopted measures mitigating these threats by following the guidelines proposed by Wohlin et al(Wohlin2012ESE, ). Internal validity is not considered, since this study does not address any causal relationships between variables.
@@ 6.1. Construct Validity
Construct validity concerns the connection between operational measures and the studied subjects. The first potential threat to construct validity is the evaluation of our selected ML/DL-based and LLM-based classifiers. Regarding the classifiers’ performance evaluation, we used the four metrics (recall, precision, F1-score, and accuracy) that are widely used to evaluate the performance of various automated software engineering techniques (see Section 7.3). Therefore, the metrics pose little threat to the construct validity. Another potential threat comes from the convenience sampling method of the survey and interview, as individuals with negative views towards the research might be less inclined to participate in our survey and interview. This is a compromise inherent to the convenience sampling method, and it implies that this threat cannot be completely mitigated. Nevertheless, to improve the response rate and collect as many opinions as possible from practitioners, we sent two rounds of reminder emails to the involved practitioners. It is also possible that some survey respondents did not understand the statements well. To reduce this threat to construct validity, we conducted a pilot survey for refining the survey design. Besides, we provided an “I Don’t Know” option in our survey. Moreover, we also interviewed six participants to collect in-depth opinions besides the six defined statements in the survey. To mitigate this threat, we offered an optional open-ended question that allows respondents to freely share their thoughts with us.
@@ 6.2. External Validity
External validity pertains to the generalizability of our experiment results and findings. In certain situations, textual artifacts (e.g., code review, architecture discussions) may be unavailable. In such scenarios, we recognize that our approach may be complemented with traditional code-based architecture conformance checking techniques. Exploring the integration of such an automated approach into existing IDEs or tools remains part of our future work. As noted by Peter et al. (peters2017text, ), it is not possible to use an exhaustive approach to test all learning models. Therefore, we selected five common ML algorithms, one DL algorithm, three popular pre-trained word embedding models, and three well-performed LLMs, as they are all widely used to conduct text classification. We admit that using different algorithms, pre-trained word embedding models, and LLMs may generate varying results, despite their popularity. We cannot completely mitigate this threat. Selecting different datasets and experimental settings might influence the generalizability of (part of) our findings. We acknowledge that the chosen open-source projects are not representative of all projects from the OpenStack and Qt platforms, and our findings may not be generalized to diverse projects, particularly those from different domains such as closed-source projects or projects on other platforms like Apache. Furthermore, the participants of the survey and interviews are contributors to the selected projects, therefore, the findings of RQ3 might not be generalized to other projects. To support further generalizability, we have made our dataset and scripts (onlinepackage, ) available to enable other researchers and practitioners to replicate, reuse, and extend our study.
@@ 6.3. Reliability
Reliability reflects the replicability of a study regarding yielding the same or similar results when other researchers reproduce this study. One potential threat to reliability comes from the experimental settings (e.g., input data and parameters for training classifiers) used in this work. To address this, we used the dataset from our previous study (Li2023wvs, ) that had publicly shared the labeled data and corresponding scripts for collecting data. Besides, we also shared the source code in our replication package (onlinepackage, ) to facilitate reproduction and validation. Another threat might concern the replicability of the survey and interview. In the current study, biases related to survey design were mitigated through careful question formulation and pilot testing. While such biases cannot be entirely eliminated, we provided the survey template, semi-structured interviews, and customized emails in our replication package (onlinepackage, ) to maximize transparency and reliability, in order to facilitate replication in future studies.
@@ 7. Related Work
In this section, we briefly discuss the current state of research in three associated areas of our study. We first describe related work on code review comments in Section 7.1; we then introduce previous studies on architecture violations in Section 7.2; we finally present the benefits of utilizing ML/DL and LLM-related techniques to analyze artifacts in software development in Section 7.3.
@@ 7.2. Architecture Violations
Besides, Terra et al(Terra2015rsr, ) reported their experience in fixing architectural violations. They proposed a recommendation system ArchFix that provides refactoring guidelines for developers and maintainers to repair architectural violations in the module architecture view of object-oriented systems. Their results show that their approach can provide correct recommendations for 75% and 79% of the detected architecture violations on two selected systems, respectively. Maffort et al(Maffort2016mav, ) proposed an approach to check architecture conformance for detecting architecture violations based on defined heuristics. They claimed that their approach relies on the defined heuristic rules and can rapidly raise architecture violation warnings.
7.3. Analyzing Software Repositories with ML/DL and LLM Techniques